# Artificial Intelligence Graduate School (AIGS) Computer Science and Engineering (CSE)

### v1

@@ 3.3 Bottleneck Reordering
Although each patch tunnel can retain its memory usage under the memory budget $m_{b}$ by adjusting the convolution kernel parameters, i.e., $s,k_{h}$, $w_{h}$, $p_{h}$, $p_{w}$, and $c_{out}$ in Eq. 2, the extremely low memory budget of many low-end devices (e.g., $m_{b}=256$ KB) constrains the number of output channels, i.e., $c_{out}$, too small to learn features sufficiently, likely to cause the model performance drop. In other words, to secure a reasonable level of model performance, $c_{out}$ should be increased sufficiently large beyond the memory budget $m_{b}$.
@@
Figure 5: The execution order of convolutions in the bottleneck, i.e., 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, is rearranged to (1, 4, 7, 8, 9), (2, 5, 10, 11, 12), (3, 6, 13, 14, 15) by ‘bottleneck reordering’, reducing the peak memory requirement of two intermediate outputs (i.e., $Y^{p}$ and $Y^{d}$) into $\frac{1}{c_{out}^{p}}$. The red and blue boxes indicate the memory requirement with and without ‘bottleneck reordering’, respectively.
@@ 4 Experiment
We implement a set of CNNs using the proposed three design principles with C language, which is available at an anonymous GitHub1. We deploy them onto STM32 development boards [25], equipped with an ARM Cortex-M7 processor and SRAM ranging between 256 and 4,096 KB, and evaluate them with two representative tasks, i.e., image classification [5, 3] and object detection [9, 19].
@@ 4.1 Image Classification
Model Name Peak-mem Quant Res Top-1 Top-5
Proposed network 254KB fp32 224 63.84% 84.80%
63KB int8 224 61.58% 83.26%
533KB fp32 224 64.68% 85.47%
133KB int8 224 61.50% 83.19%
MCUNetV1 968KB fp32 160 60.90% 83.30%
242KB int8 160 60.30% 82.60%
MCUNetV2 196KB int8 N/A 64.90% 86.20%
456KB int8 N/A 71.80% 90.70%
MobileNetV2 5,619KB fp32 224 72.83% 91.06%
Model Name Peak-mem Quant Res Accuracy
Proposed network 70KB fp32 66 82.80%
18KB int8 66 82.63%
254KB fp32 224 90.67%
63KB int8 224 87.99%
MCUNetV1 584KB fp32 64 87.40%
146KB int8 80 87.30%
648KB fp32 160 88.90%
162KB int8 160 88.90%
1,244KB fp32 144 91.70%
311KB int8 144 91.80%
MCUNetV2 30KB int8 N/A 90.00%
62KB int8 N/A 93.00%
118KB int8 N/A 94.00%
MobileNetV2 5,619KB fp32 224 91.17%
Channel Output. Tab. 3 shows the peak memory usage and classification accuracy of ImageNet [5] with different numbers of output channels, i.e., $c_{out}$ in Eq. 2, of the point-wise (expansion) convolution in the bottleneck. As shown in the table, the peak memory usage does not increase over $c_{out}$, enabled by ‘bottleneck reordering’, whereas the classification accuracy increases, e.g., 58.82% ($c_{out}{=}2$) and 63.84% ($c_{out}{=}8$) top-1 accuracy, entailing a slight increase in the point-wise kernel size. It shows that a flexible trade-off exists between the model performance and kernel size, whereas the peak memory usage remains constant.
Table 3: The peak memory usage and classification accuracy on ImageNet [5] with various output channels (i.e., $c_{out}$) of the point-wise (expansion) conv in the bottleneck.
   $\mathbf{c_{out}}$ Peak-mem Quant Top-1 Top-5
2 254KB fp32 58.82% 81.04%
2 63KB int8 55.75% 78.97%
4 254KB fp32 62.28% 83.76%
4 63KB int8 58.98% 81.63%
6 254KB fp32 63.48% 84.98%
6 63KB int8 60.70% 82.92%
8 254KB fp32 63.84% 84.80%
8 63KB int8 61.58% 83.26%
Patch Composition. Tab. 4 shows the peak memory usage and classification accuracy of ImageNet [5] and VWW [3] when removing the central patch tunnel from the network. As shown in the table, the classification accuracy drops for both ImageNet [5] (from 63.84% to 62.32%) and VWW [3] (from 82.80% to 78.97%), substantiating the crucial role of the central patch in classification.
@@
Dataset # of Patches Peak-mem Quant Top-1 Top-5
ImageNet 4 254KB fp32 62.32% 84.38%
ImageNet 4 63KB int8 61.11% 83.17%
VWW 4 70KB fp32 78.97% N/A
VWW 4 18KB int8 78.46% N/A
LABEL:{tab:experiment_different_patches} shows the peak memory usage and classification accuracy of ImageNet [5] and VWW [3] over different numbers of patches (i.e., from $k=5$ to $k=2$), which learns the non-segmented entire images resized to $[130{\times}130{\times}3]$ and $[42{\times}42{\times}3]$, respectively, where the two top rows show the cases when learning the proposed five segmented image patches (i.e., ‘input segmentation’). It shows that ‘input segmentation’ helps boost the model performance, while learning the same entire images as multiple patches does not provide similar performance regardless of the number of patches (e.g., 63.84% vs. 60.70% in ImageNet [5]), demonstrating the proposed ‘input segmentation’ is essential to achieve competitive performance.
@@ 4.2 Object Detection
Memory Usage and Model Performance. Tab. 7 and 8 show the proposed networks with different memory budgets and mAP (mean Average Precision) on PASCAL-VOC [9] and MS-COCO [19], compared with MobileNet [23] and MCUNet [17, 16]. For MS-COCO[19], all networks are trained with trainval35k and evaluated on minival5k in the MS-COCO 2014 dataset. Similar to image classification, the proposed network takes the lowest memory among all in Tab. 7, i.e., 65 KB with quantization (int8), 155x and 3.8x smaller than MobileNet (10,080 KB) and MCUNet (247 KB), respectively, achieving 50.7% mAP, showing that it effectively trades off the peak memory usage and mAP. However, the network taking the second lowest memory, i.e., 176 KB, achieves 62.1% mAP, comparable to MCUNet taking 247 KB to provide 64.6% mAP. In Tab. 8, the proposed network also takes much smaller memory compared to MobileNet, e.g., 1,719 KB (5.8x smaller) with a reasonable level of mAP (30.0%), verifying that the proposed design principles also effectively apply to object detection. Fig. 6 shows examples of object detection on PASCAL-VOC [9] and MS-COCO [19].
@@
Model Name Peak-mem Quant Res $\mathbf{mAP}$
Proposed network 259KB fp32 130 51.8%
65KB int8 130 50.7%
702KB fp32 224 63.4%
176KB int8 224 62.1%
1,260KB fp32 300 67.0%
315KB int8 300 66.3%
MCUNetV1 466KB int8 224 51.4%
MCUNetV2 438KB int8 224 68.3%
247KB int8 224 64.6%
MobileNetV2 10,080KB fp32 300 68.6%
@@
Model Name Peak-mem Quant Res $\mathbf{mAP^{val}}$
Proposed network 446KB fp32 130 22.0%
112KB int8 130 21.2%
1,210KB fp32 224 28.9%
303KB int8 224 27.2%
1,719KB fp32 224 30.0%
430KB int8 224 28.0%
MobileNetV2 10,080KB fp32 300 31.8%
@@
Channel Output. Tab. 9 shows the peak memory usage and mAP (mean Average Precision) on PASCAL-VOC [9] and MS-COCO [19] with different numbers of output channels, i.e., $c_{out}$ in Eq. 2, of the point-wise (expansion) convolution in the bottleneck. As shown in the table, the peak memory usage does not increase over $c_{out}$, enabled by ‘bottleneck reordering’, whereas mAP increase, e.g., from 52.3% to 63.4% in PASCAL-VOC [9], along with a slight increase in the convolution kernel size. Similar to the image classification tasks, the proposed network provides a flexible trade-off between mAP and the size of the kernel for object detection, while keeping the peak memory requirement constant.
Table 9: The object detection on PASCAL-VOC [9] and MS-COCO [19] with different output channels (i.e., $c_{out}$) of the point-wise (expansion) conv within the bottleneck.
Pascal-VOC $\mathbf{c_{out}}$ Mem Quant $\mathbf{mAP}$ 2 702KB fp32 52.3% 2 176KB int8 46.6% 4 702KB fp32 57.3% 4 176KB int8 53.1% 6 702KB fp32 61.5% 6 176KB int8 60.2% 8 702KB fp32 63.4% 8 176KB int8 62.1% MS-COCO $\mathbf{c_{out}}$ Mem Quant $\mathbf{mAP^{val}}$ 2 1,719KB fp32 22.4% 2 430KB int8 13.3% 4 1,719KB fp32 26.6% 4 430KB int8 22.9% 6 1,719KB fp32 28.3% 6 430KB int8 25.5% 8 1,719KB fp32 30.0% 8 430KB int8 28.0%
@@ 5 Limitations and Discussions
5.1 Network Size
Although the proposed network takes much smaller memory when compared to state-of-the-art memory-efficient MobileNet [23] and MCUNet [17, 16], e.g., 89x and 3.1x smaller for ImageNet [5], respectively, the network size is not decreased significantly, unlike huge memory reduction. That is because the number of weight parameters of the proposed network for ImageNet [5], i.e., 3.34 million, is similar to that of MobileNet (i.e. 3.2 million) and larger than that of MCUNet (i.e. 0.73 million). It can be regarded as a trade-off between memory usage and network size, which can be flexibly adjusted based on the system and/or application requirements. Considering that the MCUNet [17, 16] architecture has been obtained by NAS (Neural Architecture Search) [6, 12, 8], unlike the proposed network built without NAS, we expect that more efficient network architecture in terms of the number of weight parameters could be identified by using NAS, while the proposed three methods are kept as basic principles of architectural design at the same time. Alternatively, the weight parameters of $k$ patch tunnels could be shared among them to reduce the total number of weight parameters of the network by $k$ times. Also, it can be considered to combine some patch tunnels at the deeper layer of the network where the memory usage is smaller, as seen in Fig. 1. We will investigate these approaches in future work.
5.2 Larger Images
The proposed ‘input segmentation’ and ‘patch tunneling’, in principle, can be applied to images larger than the ImageNet data (i.e. $224\times 224$[5] by adjusting the number of patches and the overlapped areas accordingly based on the memory budget. However, as a large image is segmented into an increased number of smaller patches, the amount of information included in a single patch tends to decrease, making it more difficult to look at and learn the image from a global perspective. As each patch tunnel separately learns features available only in an individual local patch, the feature aggregation procedure for the global image learning performed at the last layer of the network is likely to become more difficult, possibly causing model performance degradation. We expect that it can be mitigated by segmenting an image into patches more flexibly, such as making different sizes of patches at various positions of the image which can cover different areas of the image with varying sizes. We also leave it for future work.
5.3 Other Tasks and Datasets
In this work, we construct memory-efficient CNNs for two vision tasks, i.e., image classification and object detection, demonstrating its effectiveness. For a more thorough evaluation, however, additional experiments on different tasks and various datasets, such as image segmentation [4, 29], super resolution [22, 7], and image generation [21, 28], should be conducted. We plan to apply the proposed methods to various tasks and investigate their applicability and limitations.
@@ 6 Conclusion
We introduce a bottleneck-based extremely memory-efficient CNN that fits the tight memory capacity (e.g., under 256 KB) of low-end embedded and IoT devices. The proposed network is constructed by the three design principles, i.e., ‘input segmentation’, ‘patch tunneling’, and ‘bottleneck reordering’, providing a flexible way of adjusting the memory usage of a CNN based on the memory budget of the low-end device. We implement the proposed network on a real low-end device (i.e., STM32H7 [25]), demonstrating that the proposed design principles can build memory-aware CNNs for various vision tasks running on memory-constrained low-end devices, e.g., enabling memory-efficient image classification on ImageNet [5] using 63 KB of memory (61.58% top-1 accuracy), and objection detection (PASCAL-VOC [9]) using 65 KB of memory (50.7% mAP).

### v2

@@ 3.3 Bottleneck Reordering
Although each patch tunnel can retain its memory usage within the memory budget $m_{b}$ by adjusting convolution kernel parameters, i.e., $s,k_{h}$, $w_{h}$, $p_{h}$, $p_{w}$, and $c_{out}$ in Eq. 2, the extremely low memory budget of many low-end devices (e.g., $m_{b}=256$ KB) limits the number of output channels i.e., $c_{out}$, to a point where feature learning becomes insufficient, potentially degrading model performance. In [23], the expansion ratio $t$ was introduced to expand $c_{out}$ in point-wise (expansion) convolution. Then $c_{out}$ can be expressed in terms of $t$, such as $c_{out}=c_{in}\times t$. For adequate model performance, however, $c_{out}$ must be increased beyond the constraints of the memory budget $m_{b}$.
@@
Figure 5: In a typical bottleneck operation, the black 1, 2, and 3 conv layers are processed sequentially. With bottleneck reordering, the layers are computed in the order of red (1, 2, 3), (4, 5, 6), and (7, 8, 9). After each cycle, two intermediate outputs are freed, and the next outputs (gray boxes and lines) are allocated. This approach reduces the peak memory requirement caused by intermediate outputs to $1/c_{\text{out}}$.
@@ 4 Experiment
We implement a series of CNNs based on the proposed three design principles in C language and deploy them onto STM32 development boards [25], featuring an ARM Cortex-M7 processor and 256 to 4,096 KB of SRAM. The networks are evaluated on two key tasks: image classification [5, 3] and object detection [9, 19].
@@
Model Name Peak-mem(KB) Quant Res Top-1(%) Top-5(%)
Proposed network 254/63 fp32/int8 224 63.84/61.58 84.80/83.26
533/133 fp32/int8 224 64.68/61.50 85.47/83.19
MCUNetV1 968/242 fp32/int8 160 60.90/60.30 83.30/82.60
MCUNetV2 196 int8 N/A 64.90 86.20
456 int8 N/A 71.80 90.70
MobileNetV2 5,619 fp32 224 72.83 91.06
Model Name Peak-mem(KB) Quant Res Accuracy(%)
Proposed network 70/ 18 fp32/ int8 66 82.80/ 82.63
254/ 63 fp32/ int8 224 90.67/ 87.99
MCUNetV1 584/ 146 fp32/ int8 64 87.40/ 87.30
648/ 162 fp32/ int8 160 88.90/ 88.90
1,244/ 311 fp32/ int8 144 91.70/ 91.80
MCUNetV2 30 int8 N/A 90.00
62 int8 N/A 93.00
118 int8 N/A 94.00
@@ 4.1 Image Classification
Channel Output. Tab. 3 shows the peak memory usage and classification accuracy of ImageNet [5] with different numbers of expansion ratio, i.e., $t$, of the point-wise (expansion) convolution in the bottleneck. As shown in the table, the peak memory usage does not increase over $t$, enabled by ‘bottleneck reordering’, whereas the classification accuracy increases, e.g., 58.82% ($t{=}2$) and 63.84% ($t{=}8$) top-1 accuracy, entailing a slight increase in the point-wise kernel size. It shows that a flexible trade-off exists between the model performance and kernel size, whereas the peak memory usage remains constant.
Table 3: The peak memory usage and classification accuracy on ImageNet [5] with various expansion ratios (i.e., $t$) in the bottleneck. The value of $c_{out}$, representing the number of output channels after channel expansion, is derived by multiplying $t$ with the number of output channels from the preceding depthwise convolution layer.
$\mathbf{t}$ Peak-mem(KB) Quant Top-1(%) Top-5(%)
2 254/ 63 fp32/ int8 58.82/ 55.75 81.04/ 78.97
4 254/ 63 fp32/ int8 62.28/ 58.98 83.76/ 81.63
6 254/ 63 fp32/ int8 63.48/ 60.70 84.98/ 82.92
8 254/ 63 fp32/ int8 63.84/ 61.58 84.80/ 83.26
Patch Composition. Tab. 4 shows the peak memory usage and classification accuracy of ImageNet [5] and VWW [3] when removing the central patch tunnel from the network. As shown in the table, the classification accuracy drops for both ImageNet (from 63.84% to 62.32%) and VWW (from 82.80% to 78.97%), substantiating the crucial role of the central patch in classification. LABEL:{tab:experiment_different_patches} shows the peak memory usage and classification accuracy of ImageNet [5] and VWW [3] over different numbers of patches (i.e., from $k=5$ to $k=2$), which learns the non-segmented entire images resized to $[130{\times}130{\times}3]$ and $[42{\times}42{\times}3]$, respectively, where the two top rows show the cases when learning the proposed five segmented image patches (i.e., ‘input segmentation’). It shows that ‘input segmentation’ helps boost the model performance, while learning the same entire images as multiple patches does not provide similar performance regardless of the number of patches (e.g., 63.84% vs. 60.70% in ImageNet [5]), demonstrating the proposed ‘input segmentation’ is essential to achieve competitive performance.
@@
Dataset # of Patches Peak-mem(KB) Quant Top-1(%) Top-5(%)
ImageNet 4 254/63 fp32/int8 62.32/84.38 84.38
VWW 4 70/18 fp32/int8 78.97/78.46 N/A
@@ 4.2 Object Detection
Memory Usage and Model Performance. Tab. 7 and 8 show the proposed networks with different memory budgets and mAP (mean Average Precision) on PASCAL-VOC [9] and MS-COCO [19], compared with MobileNet [23] and MCUNet [17, 16]. For MS-COCO, all networks are trained with trainval35k and evaluated on minival5k in the MS-COCO 2014 dataset. Similar to image classification, the proposed network takes the lowest memory among all in Tab. 7, i.e., 65 KB with quantization (int8), 155x and 3.8x smaller than MobileNet (10,080 KB) and MCUNet (247 KB), respectively, achieving 50.7% mAP, showing that it effectively trades off the peak memory usage and mAP. However, the network taking the second lowest memory, i.e., 176 KB, achieves 62.1% mAP, comparable to MCUNet taking 247 KB to provide 64.6% mAP. In Tab. 8, the proposed network also takes much smaller memory compared to MobileNet, e.g., 1,719 KB (5.8x smaller) with a reasonable level of mAP (30.0%), verifying that the proposed design principles also effectively apply to object detection. Fig. 6 shows examples of object detection on PASCAL-VOC [9] and MS-COCO.
@@
Model Name Peak-mem(KB) Quant Res mAP (%)
Proposed network 259/ 65 fp32/ int8 130 51.8/ 50.7
702/ 176 fp32/ int8 224 63.4/ 62.1
1,260/ 315 fp32/ int8 300 67.0/ 66.3
MCUNetV1 466 int8 224 51.4
MCUNetV2 438 int8 224 68.3
247 int8 224 64.6
MobileNetV2 10,080 fp32 300 68.6
@@
Model Name Peak-mem(KB) Quant Res $\mathbf{mAP^{val}}$(%)
Proposed network 446/ 112 fp32/ int8 130 22.0/ 21.2
1,210/ 303 fp32/ int8 224 28.9/ 27.2
1,719/ 430 fp32/ int8 224 30.0/ 28.0
MobileNetV2 10,080 fp32 300 31.8
Channel Output. Tab. 9 shows the peak memory usage and mAP (mean Average Precision) on PASCAL-VOC and MS-COCO with different numbers of expansion ratio, i.e., $t$, of the point-wise (expansion) convolution in the bottleneck. As shown in the table, the peak memory usage does not increase over $t$, enabled by ‘bottleneck reordering’, whereas mAP increase, e.g., from 52.3% to 63.4% in PASCAL-VOC, along with a slight increase in the convolution kernel size. Similar to the image classification tasks, the proposed network provides a flexible trade-off between mAP and the size of the kernel for object detection, while keeping the peak memory requirement constant.
Table 9: The object detection on PASCAL-VOC [9] and MS-COCO [19] with different expansion ratios (i.e., $t$) in the point-wise (expansion) conv within the bottleneck.
Pascal-VOC MS-COCO
$\mathbf{t}$ Mem(KB) Quant Res mAP(%) Mem(KB) Quant Res mAP(%)
2 446/ 112 fp32/ int8 130 22.0/ 21.2 446/ 112 fp32/ int8 130 22.0/ 21.2
2 210/ 303 fp32/ int8 224 28.9/ 27.2 446/ 112 fp32/ int8 130 22.0/ 21.2
2 719/ 430 fp32/ int8 224 30.0/ 28.0 446/ 112 fp32/ int8 130 22.0/ 21.2
@@ 5 Limitations and Discussions
Network Size. Although the proposed network takes much smaller memory when compared to state-of-the-art memory-efficient MCUNet [17, 16] and MobileNet [23], e.g., 89x and 3.1x smaller for ImageNet [5], respectively, the network size is not decreased significantly, unlike huge memory reduction. That is because the number of weight parameters of the proposed network for ImageNet, i.e., 3.34 million, is similar to that of MobileNet (i.e. 3.2 million) and larger than that of MCUNet (i.e. 0.73 million). It can be regarded as a trade-off between memory usage and network size, which can be flexibly adjusted based on the system and/or application requirements. Since the MCUNet architecture was designed using NAS, unlike our network, we expect that NAS could identify a more efficient architecture in terms of weight parameters while maintaining our three proposed design principles. Alternatively, the weight parameters of $k$ patch tunnels could be shared among them to reduce the total number of weight parameters of the network by $k$ times. Also, it can be considered to combine some patch tunnels at the deeper layer of the network where the memory usage is smaller, as seen in Fig. 1. We will investigate these approaches in future work.
Larger Images. The proposed ‘input segmentation’ and ‘patch tunneling’, in principle, can be applied to images larger than the ImageNet data (i.e. $224\times 224$) by adjusting the number of patches and the overlapped areas accordingly based on the memory budget. However, as a large image is segmented into an increased number of smaller patches, the amount of information included in a single patch tends to decrease, making it more difficult to look at and learn the image from a global perspective. As each patch tunnel separately learns features available only in an individual local patch, the feature aggregation procedure for the global image learning performed at the last layer of the network is likely to become more difficult, possibly causing model performance degradation. We expect that it can be mitigated by segmenting an image into patches more flexibly, such as making different sizes of patches at various positions of the image which can cover different areas of the image with varying sizes. We also leave it for future work.
Other Tasks and Datasets: In this work, we construct memory-efficient CNNs for two vision tasks, i.e., image classification and object detection, demonstrating its effectiveness. For a more thorough evaluation, however, additional experiments on different tasks and various datasets, such as image segmentation [4, 29], super resolution [22, 7], and image generation [21, 28], should be conducted.
@@ 6 Conclusion
We introduce a bottleneck-based, extremely memory-efficient CNN that fits the tight memory constraints (e.g., under 256 KB) of embedded and IoT devices. The network is built on three design principles: input segmentation’, patch tunneling’, and ‘bottleneck reordering’, allowing flexible adjustment of memory usage based on device capacity. We implement the network on an actual device (STM32H7 [25]), demonstrating that these design principles enable memory-aware CNNs for various vision tasks, such as memory-efficient image classification on ImageNet [5] using 63 KB (61.58% top-1 accuracy) and object detection on PASCAL-VOC [9] using 65 KB (50.7% mAP).
Acknowledgement
This research was supported by the Challengeable Future Defense Technology Research and Development Program through the Agency for Defense Development(ADD) funded by the Defense Acquisition Program Administration(DAPA) in 2022(No.915062201), and Institute of Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No.RS-2020-II201336, Artificial Intelligence Graduate School Program(UNIST)).